Metalfab Pump and Filter Inc. estimates that the cost of steel bodies for pressure valves will increase by $2 every 3 months. If the cost for the first quarter is expected to be $80, what is the present worth of the costs for a 5-year time period at an interest rate of 1.86% per quarter?

1
Expert's answer
2022-02-06T14:22:20-0500

Number of months

"4\times5=20" months

Cost taking into account the increase:

"80+2\times20=80+40=120"

Current value:

"PV=\frac{FV}{(1+i)^{20}}"


"PV=\frac{120}{(1+0.0186)^{20}}=83.01"